OpenStudy (anonymous):

Cd=3x-5, bc=2x+3, gh=7x-17 , hI=5x-12 , what is x?

OpenStudy (anonymous):

set them all equal to 180

OpenStudy (anonymous):

But what's the answer

OpenStudy (anonymous):

c=3x−5d b=2x+3c g=7x−17h

OpenStudy (anonymous):

c=−5d+3x b=3c+2x g=−17h+7x

OpenStudy (anonymous):

and thats as far as you can go

OpenStudy (anonymous):

sorry messed up now you got to plug it in the formula that idk

OpenStudy (anonymous):

nevermind that is not right